OpenStudy (anonymous):

if z=a-ai and (2z+3)^2012=(z+5)^2012 find a

OpenStudy (anonymous):

i think a = -2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

hold on i am changed my mind

OpenStudy (anonymous):

\[2z+3=(2a+3)-2ai\] \[z+5=(a+5)-ai\] when you raise these numbers two that huge power the absolute value is raised, and the angle is multiplied, so the absolute values must have been the same to begin with, in other words \[(2a+3)^2+4a^2=(a+5)^2+a^2\] and solve that quadratic. you do not get -2, i made a mistake

OpenStudy (anonymous):

thank you soooooo much :)